# Operational reconstruction of Feynman rules for quantum amplitudes via composition algebras

### v1

@@ Abstract
This paper revisits an operational model presented in [GKS2010] as part of the Quantum Reconstruction Program, describing transition amplitudes between measurements. Our methodology establishes clarity by separating axioms from mathematics, choices from physics, and deductions therefrom. We carefully evaluate the original model in a coordinate-independent way without requiring a two-dimensional space a priori. All scalar field and vector space axioms are traced from model axioms and observer choices, including additive and multiplicative units and inverses. Known theorems in math classify allowable amplitude algebras as the real associative composition algebras, namely, the two-dimensional (split-)complex numbers and the four-dimensional (split-)quaternions. Observed probabilities are quadratic in amplitudes, akin to the Born rule in physics. We point out select ramifications of postulated model axioms and ways to rephrase observer questions; and advertise broad utility of our work towards follow-on discovery, whether as a consequence, generalization, or alternative. One seemingly minute generalization is sketched in the outlook, with algebraic consequences at the heart of current open questions in mathematics and physics.
@@ 2.3.2 Chaining is associative but partially noncommutative
In the other case, when both commuted expressions $\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}$ and $\mathcal{M}^{\mathrm{B}}\cdot\mathcal{M}^{\mathrm{A}}$ are defined, there is still no a priori basis for this symmetry. We don’t need to assume that changing the order of measurements, or changing the order of measurement outcomes never affects the observation.
@@ 2.3.7 Insertion
When inserting sequences of measurements and results, we use a similar notation as for the insertion of individual ones. Inserting a cyclic measurement sequence $^{\circ}\negthinspace\mathcal{M}^{\Lambda}$ after the sequence $\mathcal{M}^{\mathrm{J}}$ of a factorization $\mathcal{M}=\prod_{\mathrm{K}=1}^{\mathrm{L}}\mathcal{M}^{\mathrm{K}}$ is written as:
@@
Chaining requires $\mathbf{T}\left(\mathcal{M}^{\mathrm{J}-1}\right)=\mathbf{S}\left({}^{\circ}\negthinspace\mathcal{M}^{\Lambda}\right)=\mathbf{T}\left({}^{\circ}\negthinspace\mathcal{M}^{\Lambda}\right)=\mathbf{S}\left(\mathcal{M}^{\mathrm{J}}\right)$ and this operation is only allowed for cyclic $^{\circ}\negthinspace\mathcal{M}^{\Lambda}$. It increases sequence length by all elements of $^{\circ}\negthinspace\mathcal{M}^{\Lambda}$ that are not its source or target:
$\ell_{\mathcal{M}\downarrow_{\mathrm{J}}{}^{\circ}\negthinspace\mathcal{M}^{\Lambda}}=\ell_{\mathcal{M}}+\ell_{{}^{\circ}\negthinspace\mathcal{M}^{\Lambda}}-2.$ (2.45)
@@ 3.1 Paths have probability and establish closure
A general path $A=\left[A_{1},A_{2},\ldots,A_{\ell}\right]\in\mathcal{S}_{\mathcal{M}}$ over measurements $\mathcal{M}=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,\mathbf{M}_{\ell}\right]\in\left[\mathbb{M}\right]$, $\ell\geq 2$, will now be associated with a probability $P\in\left[0,1\right]\in\mathbb{R}$. Each measurement $\mathbf{M}_{j}$ has $\left|\mathbf{M}_{j}\right|$ detectors and a total of $\prod_{j=1}^{\ell}\left|\mathbf{M}_{j}\right|$ different paths can be observed. From repeated experiments using the same setup, a table can then be compiled that counts how often each allowable path occurred.
@@ 3.2 Strong repeatability and factorization
Paths must start and end with atomic measurements and have atomic source and target outcomes. If a measurement sequence $\mathcal{M}^{\mathrm{A}}$ contains another atomic measurement, $^{\alpha}\mathbf{M}_{j}$ with $j\neq 1,\ell$, then we can decompose $\mathcal{M}^{\mathrm{A}}$ into two sequences, $\mathcal{M}^{\mathrm{B}}$ and $\mathcal{M}^{\mathrm{C}}$, and understand that process as effectively duplicating $^{\alpha}\mathbf{M}_{j}$ and designating it as $\mathbf{T}\left(B\right)$ and $\mathbf{S}\left(C\right)$, respectively:
@@
This process of unchaining can be repeated on $\mathcal{M}^{\mathrm{A}}$ until all resulting measurements $\mathcal{M}^{\mathrm{J}}$, $\mathrm{J}=1\ldots\mathrm{L}$, and paths $A^{\mathrm{J}}$ are atomic only at their respective source and target:
@@ 3.6 Impossible paths
Coarsening may only modify exactly one element in a measurement sequence. There are two cases: (1) the sequence step modified under coarsening is not part of an IGP, or (2) the sequence step is one of an IGP. In case (1), the resulting path must also be impossible, since no IGP is modified. In the case of (2), only one of the two IGPs is modified. If the result would make the IGP possible, this would mean that the resulting path contains the same underlying atomic outcome element in both measurement results of the original IGP. This would be a contradiction because only one of the results is modified, and both initial paths are impossible and therefore do not contain any shared underlying atomic outcome elements at the IGP. A similar argument applies to refinement. ∎
@@ 5.2 Quadratic probabilities
Conversely, the probability of independently observing any of the paths $\left[N,M^{r}\right]$ must also be $1$,
@@
The sum of quadratic forms as in (5.13) is again a quadratic form, and the space of signomials of the same degree is closed under summation likewise (5.14). For both these sums to always add up to $1$ over the same coefficients from all $\mathbf{a}_{r}$, the exponential behavior of all argument components must be identical. This identifies $\alpha=2$ and the $p$-functions with $Q$,
@@ 5.6 A note on split-algebras
Compared to the split-algebras with their hyperbolic $p$-function, the circular geometry in $\mathbb{C}$ and $\mathbb{H}$ may appear simpler or more symmetric when separating formally defined abstract amplitudes from those with $p\in\left[0,1\right]$ that may represent concrete paths in the model. However, this does not exclude split-algebras from utility towards reconstructing Quantum Mechanics: The experiment ultimately verifies nature’s geometry, regardless of how simple or symmetric we may find it. Consequently, in this paper, we refrain from making further assumptions or postulates that have no effect other than arriving at a more desirable or compact result.
@@ 6.1 What has and has not been accomplished
In this paper, we followed up on a model that was first advertised in [GKS2010] as a building block in the Quantum Reconstruction Program, which included reconstructing core algebraic properties brought initially forward by Richard P. Feynman. Addressing open questions posed in [Koepl2023gks], we confirmed that it is possible to represent probabilities from the model by amplitudes on complex numbers. Going beyond the original work, we did so coordinate-independently without requiring a two-dimensional configuration space a priori. Furthermore, we traced algebraic unit elements, scalar multiplication, and most field axioms for scalars from the model. Demonstrating that scalars form an algebraic field, including invertibility of addition and multiplication, allowed us to conclude that the configuration space of amplitudes is indeed a vector space [^18] We believe this to be a significant innovation, also with thanks to John A. Shuster for pointing out a counterexample early on: If one were to define multiplication $\odot$ and addition $\oplus$ between a pair of two-dimensional vectors $\mathbf{a}$ and $\mathbf{b}$ in terms of familiar complex number operations, $\mathbf{a}\odot\mathbf{b}:=\exp\left(\ln\mathbf{a}\ln\mathbf{b}\right)$, $\mathbf{a}\oplus\mathbf{b}:=\mathbf{a}\mathbf{b}$, then such a system would satisfy many axioms of a conventional algebra over a field, including distributivity of multiplication over addition; however, scalar multiplication would not be homogeneous since for example $\left(\lambda\mathbf{a}\right)\odot\mathbf{b}=\exp\left(\ln\left(\lambda\mathbf{a}\right)\ln\mathbf{b}\right)=\exp\left(\ln\lambda\ln\mathbf{b}\right)\left(\mathbf{a}\odot\mathbf{b}\right)\neq\lambda\left(\mathbf{a}\odot\mathbf{b}\right)$. It would be consistent with all axioms stated in the original work but not support the conclusion. In general, algebraic constructs that result from dropping homogeneity of vector multiplication, $\left(\lambda\mathbf{a}\right)\odot\left(\mu\mathbf{b}\right)=\lambda\mu\left(\mathbf{a}\odot\mathbf{b}\right)$, while maintaining distributivity of multiplication over addition, have been investigated under the “commutative hyperoperation” label and proposed as early as [Bennett1915].. We showed that probabilities are quadratic in amplitudes without needing to know the dimensionality of the vector space. All along, we followed a program that aimed at providing clarity of thought at the interface of mathematics and physics: We separated axioms from choices in an attempt to differentiate what properties of the model derive from foundational assumptions, in contrast to properties that trace to choices made by an outside experimenter in order to provide answers for specific questions.
@@
From the physics point of view, we hope that our improved clarity in treating this elementary building block of quantum reality provides insight into existing follow-on work that considers composite systems [Goyal2014] and distinguishability [Goyal2015] and allows the model to be incorporated into new or existing work that reconstructs spacetime, the light cone, and the dynamical interaction of quantum systems [^19] Prominent graph-based approaches that address these questions include Causal Set theory, which offers powerful insights into proposed structures of spacetime, and the intriguing spin networks of Loop Quantum Gravity. For recent review articles, see e.g. [Rovelli2012, Surya2019, AB2021]..
@@ 6.2 Histories matter for amplitudes?
Algebraically, this permits algebras, where multiplication is not necessarily associative. The reader readily recognizes the nonassociative octonions $\mathbb{O}$ and split-octonions $\mathbb{O}^{\prime}$ to become permissible [Baez2002] since associativity of multiplication was the only property that ruled out these real composition algebras in this paper. Just as for the complexes and quaternions, their $p$-function models probability, which does remain invariant when modifying product association. This, therefore, still satisfies the weak closure requirement for probabilities.
While algebraic nonassociativity may feel inconvenient at first, it is essentially no different than noncommutativity in the quaternions $\mathbb{H}$, which likewise does not affect the $p$-function of the product [^20] Or noncommutativity in matrix multiplication, which does not affect the determinant.. Even more so, building an algebra from one-sided multiplication only and letting products act on a separate element $\psi$ of that algebra multiplicatively forms an associative algebra by definition: It simply forces all brackets a certain way. Using such one-sided multiplication algebras built from octonions with real coefficients, $\mathbb{O}$, with complex coefficients $\mathbb{C}\otimes\mathbb{O}$, quaternionic coefficients $\mathbb{H}\otimes\mathbb{O}$, and iteratively further “tensoring” these division algebras to form algebras like $\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$ and larger, all form Clifford algebras; and that separate element $\psi$ these act on multiplicatively hence is identified as a spinor [^21] Again, by definition, a spinor can be understood as an element on which a Clifford algebra acts.. This construction has been explored in physics for use in modeling symmetries in the Standard Model of particles since the late 20th century [Dixon1994, Dixon2014]. Significant recent progress towards a deeper, more complete understanding of the various algebraic aspects, possible realizations, degrees of freedom, consequences, and constraints [Furey2016, FuH2022-OneGen, FuH2022-SymmBr, Furey2023rmap1] led to active research in the field. Curiously, these endeavors use the same algebras that would become permissible in our model here when allowing weak closure for answering an experimenter’s question.
@@ 6.3 Do all histories matter?
The occurrence of related algebraic structures in different contexts, now even tracing from different model assumptions, leads one to wonder about a common provenance. How could our current “histories don’t matter” postulate be reconcilable with an “all histories matter” axiom? Closing on an inspirational note, when weakening closure from “histories don’t matter” to include “… for probabilities only,” to become the anthropocentric consequence of an observer approaching nature with the very question for probabilities, it is conceivable that our model here describes the behavior of quantum nature under observation [^23] Quoting [Goyal2012]: Quantum Mechanics may be “[not] a description of reality in itself [but] a description of reality as experienced by an agent”.; whereas “all histories matter” approaches like superdeterminism describe the inner workings of nature when left unobserved [^24] For a sketch of an agency-free model of Quantum Field Theory, on algebras permissible when extending our current work, see e.g.  [Furey2023osmu]..

### v2

@@ Abstract
This article explores an operational model for transition amplitudes between measurements proposed by Goyal et al. within the quantum reconstruction program. To classify suitable amplitude algebras, we distinguish mathematical axioms, physical choices, and their consequences. This leads to several improvements on the published work: Our coordinate-independent approach requires no two-dimensional amplitudes a priori. All scalar field and vector space axioms are traced from model axioms and observer choices, including additive and multiplicative units and inverses. Existing mathematical characterizations identify allowable amplitude algebras as the real associative composition algebras, namely the complex numbers and the quaternions, as well as their split forms. Observed probabilities are quadratic in amplitudes, akin to the Born rule. We examine selected implications of the proposed axioms, reformulate observer questions, and highlight the broad applicability of our framework to subsequent discovery.
@@ 2.3.2 Chaining is associative but partially noncommutative
In the other case, when both commuted expressions $\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}$ and $\mathcal{M}^{\mathrm{B}}\cdot\mathcal{M}^{\mathrm{A}}$ are defined, there is still no a priori basis for this symmetry. We don’t need to assume that changing the order of measurements or changing the order of measurement outcomes never affects the observation.
@@ 2.3.7 Insertion
When inserting sequences of measurements and results, we use a similar notation as for the insertion of individual ones. Inserting a cyclic measurement sequence $^{\circ}\negthinspace\mathcal{M}^{\Lambda}$ after the sequence $\mathcal{M}^{\mathrm{J}}$ of a sequence chain $\mathcal{M}^{\mathrm{1}}\cdot\ldots\cdot\mathcal{M}^{\mathrm{L}}$ is written as:
@@
Chaining requires $\mathbf{T}\left(\mathcal{M}^{\mathrm{J}-1}\right)=\mathbf{S}\left({}^{\circ}\negthinspace\mathcal{M}^{\Lambda}\right)=\mathbf{T}\left({}^{\circ}\negthinspace\mathcal{M}^{\Lambda}\right)=\mathbf{S}\left(\mathcal{M}^{\mathrm{J}}\right)$ and this operation is only allowed for cyclic $^{\circ}\negthinspace\mathcal{M}^{\Lambda}$. It increases sequence length by all elements of $^{\circ}\negthinspace\mathcal{M}^{\Lambda}$ that are not its source (or target):
$\ell_{\mathcal{M}\downarrow_{\mathrm{J}}{}^{\circ}\negthinspace\mathcal{M}^{\Lambda}}=\ell_{\mathcal{M}}+\ell_{{}^{\circ}\negthinspace\mathcal{M}^{\Lambda}}-1.$ (2.45)
@@ 3.1 Paths have probability and establish closure
A general path $A=\left[A_{1},A_{2},\ldots,A_{\ell}\right]\in\mathcal{S}_{\mathcal{M}}$ over measurements $\mathcal{M}=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,\mathbf{M}_{\ell}\right]\in\left[\mathbb{M}\right]$, $\ell\geq 2$, will now be associated with a probability $P\in\left[0,1\right]\in\mathbb{R}$. Each measurement $\mathbf{M}_{j}$ has $\left|\mathbf{M}_{j}\right|$ detectors and a total of $\prod_{j=1}^{\ell}\left|\mathbf{M}_{j}\right|$ different paths can be observed over $\mathcal{S}_{\mathcal{M}}$. From repeated experiments using the same setup, a table can then be compiled that counts how often each allowable path occurred.
@@ 3.2 Strong repeatability and factorization
Paths must start and end with atomic measurements and have atomic source and target outcomes. If a measurement sequence $\mathcal{M}^{\mathrm{A}}$ contains another atomic measurement, $^{\alpha}\mathbf{M}_{j}$ with $j\neq 1,\ell$, then we can decompose $\mathcal{M}^{\mathrm{A}}$ into two sequences, $\mathcal{M}^{\mathrm{B}}$ and $\mathcal{M}^{\mathrm{C}}$, effectively duplicating $^{\alpha}\mathbf{M}_{j}$ and designating it as $\mathbf{T}\left(B\right)$ and $\mathbf{S}\left(C\right)$, respectively:
@@
This process of unchaining can be repeated on $\mathcal{M}^{\mathrm{A}}$ until all measurements $\mathcal{M}^{\mathrm{J}}$, $\mathrm{J}=1\ldots\mathrm{L}$, and paths $A^{\mathrm{J}}$ are atomic only at their respective source and target:
@@ 3.6 Impossible paths
Coarsening may only modify exactly one element in a measurement sequence. There are two cases: (1) the sequence step modified under coarsening is not part of an IGP, or (2) the sequence step is one of an IGP. In case (1), the resulting path must also be impossible, since no IGP is modified. In the case of (2), only one of the two measurements in the IGP is modified. If the result would make the IGP possible, this would mean that the resulting path contains the same underlying atomic outcome element in both measurement results of the original IGP. This would be a contradiction because only one of the results is modified, and both initial paths are impossible and therefore do not contain any shared underlying atomic outcome elements at the IGP. A similar argument applies to refinement. ∎
@@ 5.2 Quadratic probabilities
Conversely, the probability of independently observing any of the paths $\left[N,M^{r}\right]$ must also be $1$ per section 3.9,
@@
The sum of quadratic forms as in (5.13) is again a quadratic form, and the space of signomials of the same degree is closed under summation likewise (5.14). For both these sums to always add up to $1$ over the same coefficients from all $\mathbf{a}_{r}$, the exponential behavior of all summands from the argument components must be identical. This identifies $\alpha=2$ and the $p$-functions with $Q$,
@@ 5.6 A note on split-algebras
Compared to the split-algebras with their hyperbolic $p$-function, the circular geometry in $\mathbb{C}$ and $\mathbb{H}$ may appear simpler or more symmetric when separating formally defined abstract amplitudes from those with $p\in\left[0,1\right]$ that may represent concrete paths in the model. However, this does not exclude split-algebras from utility towards reconstructing Quantum Mechanics: The experiment ultimately verifies nature’s geometry, regardless of how simple or symmetric we may find it.
@@ 6.1 What has and has not been accomplished
In this paper, we explored a model that was first advertised in [GKS2010] as a building block in the Quantum Reconstruction Program, which included reconstructing core algebraic properties brought initially forward by Richard P. Feynman. Addressing open questions posed in [Koepl2023gks], we confirmed that it is possible to represent probabilities from the model by amplitudes on complex numbers. Going beyond the original work, we did so coordinate-independently without requiring a two-dimensional configuration space a priori. Furthermore, we traced algebraic unit elements, scalar multiplication, and most field axioms for scalars from the model. Demonstrating that scalars form an algebraic field, including invertibility of addition and multiplication, allowed us to conclude that the configuration space of amplitudes is indeed a vector space [^18] We believe this to be a significant innovation, also with thanks to John A. Shuster for pointing out a counterexample early on: If one were to define multiplication $\odot$ and addition $\oplus$ between a pair of two-dimensional vectors $\mathbf{a}$ and $\mathbf{b}$ in terms of familiar complex number operations, $\mathbf{a}\odot\mathbf{b}:=\exp\left(\ln\mathbf{a}\ln\mathbf{b}\right)$, $\mathbf{a}\oplus\mathbf{b}:=\mathbf{a}\mathbf{b}$, then such a system would satisfy many axioms of a conventional algebra over a field, including distributivity of multiplication over addition; however, scalar multiplication would not be homogeneous since for example $\left(\lambda\mathbf{a}\right)\odot\mathbf{b}=\exp\left(\ln\left(\lambda\mathbf{a}\right)\ln\mathbf{b}\right)=\exp\left(\ln\lambda\ln\mathbf{b}\right)\left(\mathbf{a}\odot\mathbf{b}\right)\neq\lambda\left(\mathbf{a}\odot\mathbf{b}\right)$. It would be consistent with axioms stated in the original work but not support the conclusion. In general, algebraic constructs that result from dropping homogeneity of vector multiplication, $\left(\lambda\mathbf{a}\right)\odot\left(\mu\mathbf{b}\right)=\lambda\mu\left(\mathbf{a}\odot\mathbf{b}\right)$, while maintaining distributivity of multiplication over addition, have been investigated under the “commutative hyperoperation” label and proposed as early as [Bennett1915].. We showed that probabilities are quadratic in amplitudes without needing to know the dimensionality of the vector space. All along, we followed a program that aimed at providing clarity of thought at the interface of mathematics and physics: We separated axioms from choices in an attempt to differentiate what properties of the model derive from foundational assumptions, in contrast to properties that trace to choices made by an outside experimenter in order to provide answers for specific questions.
@@
From the physics point of view, we hope that our improved clarity in treating this elementary building block of quantum reality provides insight into existing follow-on work that considers composite systems [Goyal2014] and distinguishability [Goyal2015] and allows the model to be incorporated into new or existing work that reconstructs spacetime, the light cone, and the dynamical interaction of quantum systems [^19] Prominent graph-based approaches that address these questions include Causal Set theory, which offers powerful insights into proposed structures of spacetime, and the intriguing spin networks of Loop Quantum Gravity. For recent review articles, see e.g. [Rovelli2012, Surya2019, AB2021, Surya2025]..
@@ 6.2 Histories matter for amplitudes?
Algebraically, this permits algebras, where multiplication is not necessarily associative. The reader readily recognizes the nonassociative octonions $\mathbb{O}$ and split-octonions $\mathbb{O}^{\prime}$ to become permissible [Okubo1995, Baez2002, ConwaySmithD2003] since associativity of multiplication was the only property that ruled out these real composition algebras in this paper. Just as for the complexes and quaternions, their $p$-function models probability, which does remain invariant when modifying product association. This, therefore, still satisfies the weak closure requirement for probabilities.
While algebraic nonassociativity may feel inconvenient at first, it is essentially no different than noncommutativity in the quaternions $\mathbb{H}$, which likewise does not affect the $p$-function of the product [^20] Or noncommutativity in matrix multiplication, which does not affect the determinant.. Even more so, building an algebra from one-sided multiplication only and letting products act on a separate element $\psi$ of that algebra multiplicatively forms an associative algebra by definition: It simply forces all brackets a certain way. Using such one-sided multiplication algebras built from octonions with real coefficients, $\mathbb{O}$, with complex coefficients $\mathbb{C}\otimes\mathbb{O}$, quaternionic coefficients $\mathbb{H}\otimes\mathbb{O}$, and iteratively further “tensoring” these to form algebras like $\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$ and larger, all form Clifford algebras; and that separate element $\psi$ these act on multiplicatively hence is identified as a spinor [^21] By definition, a spinor can be understood as an element on which a Clifford algebra acts.. This construction has been explored in physics for use in modeling symmetries in the Standard Model of particles since the late 20th century [Dixon1994, Dixon2014]. Significant recent progress towards a deeper, more complete understanding of the various algebraic aspects, possible realizations, degrees of freedom, consequences, and constraints [Furey2016, FuH2022-OneGen, FuH2022-SymmBr, Furey2023rmap1] led to active research in the field. Curiously, these endeavors use the same algebras that would become permissible in our model here when allowing weak closure for answering an experimenter’s question.
@@ 6.3 Do all histories matter?
The occurrence of related algebraic structures in different contexts, now even tracing from different model assumptions, leads one to wonder about a common provenance. How could our current “histories don’t matter” postulate be reconcilable with an “all histories matter” axiom? Closing on an inspirational note, when weakening closure from “histories don’t matter” to include “… for probabilities only,” to become the anthropocentric consequence of an observer approaching nature with the very question for probabilities, it is conceivable that our model here describes the behavior of quantum nature under observation [^23] Quoting [Goyal2012]: Quantum Mechanics may be “[not] a description of reality in itself [but] a description of reality as experienced by an agent”.; whereas “all histories matter” approaches like superdeterminism[AHHP2024] describe the inner workings of nature when left unobserved [^24] For a sketch of an agency-free model of Quantum Field Theory, on algebras permissible when extending our current work, see e.g. [Furey2023osmu]..